OpenStudy (anonymous):

Find the derivative of the following function -3x^2+6x-8 using the limiting process. I do not know were I am going wrong I have been getting -3deltax-6x.

OpenStudy (anonymous):

hmm why don't you show the steps you've done and I can try to check your work?

OpenStudy (anonymous):

Answer should be -6x+6

OpenStudy (anonymous):

\[-3(x+\Delta x)^2+6(x+\Delta x)-8-(-3x^2+6x-8)/\Delta x= -3(x+\Delta x)(x+\Delta x)+6x+6 \Delta x-8+3x^2-6x+8/\Delta x\]\[=-3(x^2+2\Delta x(x)+\Delta x^2+6x+6\Delta x-8+3x^2-6x+8/\Delta x=-3x^2-6x-6\Delta x-3\Delta x^2+6x+6\Delta x-8+3x^2-6x+8/\Delta x\]\[=-3\Delta x^2-6x/\Delta x=-3\Delta x-6x\]

OpenStudy (anonymous):

how do you get -6x+6?

OpenStudy (anonymous):

Look at you first part of your solution what does (x+delta x) ^ 2 expand to?

OpenStudy (anonymous):

(x+delta x)(x+delta x)

OpenStudy (kirbykirby):

Yes derivative should be -6x + 6

OpenStudy (anonymous):

x^2+2(delta x)(x) +delta (x)^2 Correct

OpenStudy (anonymous):

so than when the -3 is distributed it would equal -3x^2-6x-6delta x-3delta x^2?

OpenStudy (anonymous):

\[-3x ^{2}-6xh-3h ^{2}+6x+6h-8-[3x ^{2}+6x-8]\]

OpenStudy (anonymous):

I use h for delta x

OpenStudy (anonymous):

Oh okay i see what i did wrong thank you :)

OpenStudy (anonymous):

Welcome